# Gaitsgory’s central functor and the Arkhipov–Bezrukavnikov equivalence in mixed characteristic

### v3

@@ 4.5. Highest weight arrows
Let $\mu\in\bar{{\mathbb{X}}}_{\bullet}$ be a dominant coweight with respect to $B$. For a $\hat{G}_{\Lambda}$-representation $V$ with a single highest weight $\mu$, we see that ${\mathscr{Z}}(V)$ is supported on the $\mu$-admissible locus ${\mathcal{A}}_{{\mathcal{I}},\mu}$, cf. [AGLR22, Theorem 6.16], which equals the union of the ${\mathcal{I}}(O)$-orbits of the translations $t(\bar{\nu})$ associated with weights $\nu$ of $V$. We are going to define a canonical map
@@
But the restriction of ${\mathscr{Z}}(V)$ to the $I$-orbit ${\mathrm{Fl}}_{{\mathcal{I}},\bar{\mu}}$ is isomorphic to the local system with value $R\Gamma({\mathrm{Fl}}_{{\mathcal{I}},\bar{\mu}},{\mathscr{Z}}(V))$. On the other hand, we know by Lemma 2.12 that ${\mathrm{Fl}}_{{\mathcal{I}},\bar{\mu}}$ coincides with the intersection ${\mathcal{A}}_{{\mathcal{I}},\mu}\cap{\mathcal{S}}_{t(\bar{\mu})}={\mathrm{Fl}}_{{\mathcal{I}},\leq\bar{\mu}}\cap{\mathcal{S}}_{t(\bar{\mu})}$. Therefore, Corollary 3.24 tells us that $j^{*}_{\bar{\mu}}{\mathscr{Z}}(V)\simeq\operatorname{Grad}_{\bar{\mu}}{\mathscr{Z}}(V)[\langle 2\rho,\mu\rangle]$ in natural fashion. In particular, we get $Rj_{\bar{\mu},*}j^{*}_{\bar{\mu}}{\mathscr{Z}}(V)\simeq\operatorname{gr}_{\bar{\mu}}{\mathscr{Z}}(V)$ and we obtain the desired highest weight arrow.
@@
By 4.22 we know that the composition $\operatorname{Grad}\circ{\mathscr{Z}}$ identifies with the restriction functor from $\mathrm{Rep}_{\Lambda}(\hat{G})$ to $\mathrm{Rep}_{\Lambda}(\hat{T}^{I})$ as a tensor functor, so it is symmetric monoidal. Indeed, the monoidal structure of $\operatorname{Sat}$ comes from the monoidality of constant terms in the generic fiber ${\operatorname{Gr}}_{G}$ which is compatible with the one in the special fiber ${\mathrm{Fl}}_{{\mathcal{I}}}$, which was used in Corollary 3.28. Finally, we just have to remark that the adjunction unit is naturally symmetric monoidal as are the isomorphisms $Rj_{\bar{\mu},*}j^{*}_{\bar{\mu}}{\mathscr{Z}}(V)\simeq\operatorname{gr}_{\bar{\mu}}{\mathscr{Z}}(V)$. ∎
@@
$\mathfrak{f}_{V}\circ\mathbf{n}_{V}=0.$ (4.34)
@@ 4.6. Mixed variant
In this subsection, we are going to upgrade our previous work to the setting of mixed sheaves. We consider a $p$-adic field $F$ with ring of integers $O$, a finite residue field $k$ of cardinality $q$, and an absolute Galois group $\Gamma$. We continue to fix a quasi-split and residually split $F$-group $G$ with a Iwahori $O$-model ${\mathcal{I}}$. In this subsection, we assume furthermore that ${\Lambda}$ is an algebraic extension of ${\mathbb{Q}}_{\ell}$ and contains a preferred choice of square-root $\sqrt{q}$.
@@
Again, we can define the mixed standard functor from ${\mathscr{D}}_{\mathrm{\acute{e}t}}([\Gamma\backslash\ast])$ towards ${\mathscr{D}}_{\mathrm{\acute{e}t}}([\Gamma\backslash{\operatorname{Hk}}_{{\mathcal{I}},\bar{k}}])$
@@
both of which preserve mixed perverse sheaves by Weil II. Lemma 3.2 and Lemma 3.3 generalize to the current setting, so that we can define the mixed Wakimoto functor ${\mathscr{I}}_{\bar{\nu}}^{\mathrm{mix}}\colon{\mathscr{D}}_{\mathrm{\acute{e}t}}([\Gamma\backslash\ast])\to{\mathscr{D}}_{\mathrm{\acute{e}t}}([\Gamma\backslash{\operatorname{Hk}}_{{\mathcal{I}},\bar{k}}])$ mapping a weighted complex $M$ of ${\Lambda}$-modules to the object representing
@@
where $\bar{\nu}_{1},\bar{\nu}_{2}\in\bar{\mathbb{X}}_{\bullet}^{+}$ run over all those elements such that $\bar{\nu}=\bar{\nu}_{1}-\bar{\nu}_{2}$. Again this sends a mixed weighted ${\Lambda}$-module to a mixed perverse sheaf, since the mixedness property is preserved under derived pushforward and pullback, whereas perversity was already verified in Lemma 3.6. We can define a notion of a mixed Wakimoto complex, as lying in the full subcategory $\mathrm{Wak}^{\mathrm{mix}}$ spanned by the essential image of ${\mathscr{I}}^{\mathrm{mix}}_{\bar{\nu}}$ on mixed $\Lambda$-modules. Similarly, we can generalize Proposition 3.23 to the mixed setting, in such a way that it allows us to determine the Wakimoto grading of such an object. The full subcategory ${\mathscr{P}}(\mathrm{Wak}^{\mathrm{mix}})$ consists of the objects in $\mathrm{Wak}^{\mathrm{mix}}$ whose gradeds are all perverse.
@@
The mixed central functor ${\mathscr{Z}}^{\mathrm{mix}}$ lands in ${\mathscr{P}}(\mathrm{Wak}^{\mathrm{mix}})$. Concretely, the Wakimoto gradeds of ${\mathscr{Z}}^{\mathrm{mix}}(V)$ are canonically isomorphic to ${\mathscr{I}}_{\bar{\nu}}^{\mathrm{mix}}(V(w_{0}\bar{\nu}))$.
@@ 5.1. Generalities on coherent sheaves
Let $R$, resp. $A$ be the ring of global section on $Y$ resp. $H$. The category of finitely generated $R$-modules (which is equivalent to $\mathrm{Coh}(Y)$) has enough projectives by considering the collection of free $R$-modules. Notice that the functor of taking $H$-invariants on $R$-modules is exact by assumption on $H$. We can deduce that the $H$-equivariant free $R$-module $V\otimes M$ with $V$ being a finite dimensional representation of $H$ is projective in $\mathrm{Coh}(X)$. This collection of projectives is enough, as each coherent sheaf on $X$ is surjected upon by the $n$-fold sum of the regular representation $A\otimes R$ for $n\gg 0$, and we can find a finite representation $V\subset A$ which completes the job, by finiteness of the underlying $R$-module of the initial coherent sheaf on $X$. ∎
@@ 6. Iwahori–Whittaker averaging
Our next task is to choose a Whittaker datum. Assume $\Lambda$ is an algebraic extension of ${\mathbb{Q}}_{\ell}(\zeta)$ where $\zeta\in\bar{\mathbb{Q}}_{\ell}$ is a choice of a primitive $p$-th root of unity. We get the Artin–Schreier étale sheaf ${\mathcal{L}}_{\mathrm{AS}}$ on ${\mathbb{G}}_{a,k}$: this is the rank $1$ direct summand of the pushforward $\pi_{*}{\Lambda}$ of the constant sheaf along the Artin–Schreier cover $\pi\colon{\mathbb{G}}_{a,k}\to{\mathbb{G}}_{a,k}$ arising as the $\zeta$-eigenspace for the Galois action of ${\mathbb{Z}}/p{\mathbb{Z}}$. It is a character sheaf in the sense of Lusztig–Yun [LY20], i.e., we have isomorphisms $m^{*}{\mathcal{L}}_{\mathrm{AS}}\simeq{\mathcal{L}}_{\mathrm{AS}}\boxtimes{\mathcal{L}}_{\mathrm{AS}}$ and $e^{*}{\mathcal{L}}_{\mathrm{AS}}\cong{\Lambda}$ with respect to the multiplication $m$ and unit $e$ of ${\mathbb{G}}_{a,k}$, that satisfy associativity constraints (this is equivalent to the corresponding $\infty$-enhancement, because ${\mathscr{P}}({\mathbb{G}}_{a,k})$ is an abelian category). Besides, it satisfies the following cohomological vanishing
@@
The orbit ${\mathrm{Fl}}^{\mathrm{op}}_{{\mathcal{I}},w}$ carries a Iwahori–Whittaker local system if and only if $w$ has minimal length in its $W_{\mathrm{fin}}$-left coset, i.e., $\ell(w_{\mathrm{fin}}w)\geq\ell(w)$ for all $w_{\mathrm{fin}}\in W_{\mathrm{fin}}$. If the latter condition holds, then the rank $1$ Iwahori–Whittaker local system on ${\mathrm{Fl}}^{\mathrm{op}}_{{\mathcal{I}},w}$ is unique up to isomorphism.
@@
and claim that its composition with the quotient map ${\mathscr{P}}({\operatorname{Hk}}_{{\mathcal{I}}})\to{\mathscr{P}}_{\mathrm{as}}({\operatorname{Hk}}_{{\mathcal{I}}})$ yields a right inverse to $\mathrm{av}^{\mathrm{as}}_{{\mathcal{I}}{\mathcal{W}}}$. First notice that $\mathrm{ind}_{{\mathcal{I}}{\mathcal{W}}}(\Xi)$ is an extension of negative shifts of $\mathrm{IC}_{w}$ for $w\in W_{\mathrm{fin}}$, with the local system $\mathrm{IC}_{1}$ appearing with multiplicity $1$ (see [AR, Lemma 6.4.8]). If $w$ is non-trivial and $\mathcal{F}\in{\mathcal{P}}({\operatorname{Hk}}_{\mathcal{I}})$, we can show that $\mathrm{IC}_{w}[n]\ast{\mathcal{F}}$ is $L^{+}{\mathcal{P}}_{s}$-equivariant for some simple reflection $s$, hence its perverse cohomology groups die under the quotient map ${\mathscr{P}}({\operatorname{Hk}}_{{\mathcal{I}}})\to{\mathscr{P}}_{\mathrm{as}}({\operatorname{Hk}}_{{\mathcal{I}}})$. If $w=1$ and $n\neq 0$, then $\mathrm{IC}_{1}[n]\ast{\mathcal{F}}$ sits in non-zero perverse degree. In total, this yields an equivalence of functors $\mathrm{ind}_{{\mathcal{I}}{\mathcal{W}}}^{\mathrm{as}}\circ\mathrm{av}_{{\mathcal{I}}{\mathcal{W}}}^{\mathrm{as}}\simeq\mathrm{id}$. This implies that $\mathrm{av}_{{\mathcal{I}}{\mathcal{W}}}^{\mathrm{as}}$ is injective on $\mathrm{Ext}$ groups. To see that $\mathrm{av}_{{\mathcal{I}}{\mathcal{W}}}^{\mathrm{as}}$ is fully faithful, we argue by induction on the length of the objects being considered: the simple case is a consequence of Lemma 6.5, implying bijectivity of simple objects along $\mathrm{av}_{{\mathcal{I}}{\mathcal{W}}}^{\mathrm{as}}$; the induction step follows from the 5-lemma and the injectivity on $\mathrm{Ext}$-groups. ∎
@@ 7. Tilting modules
lands on the full subcategory of tilting objects. This is related to the ’parabolic-singular’ Koszul duality phenomenon studied by Beilinson–Ginzburg–Soergel [BGS96] for finite flag varieties and Bezrukavnikov–Yun [BY13] for Kac–Moody flag varieties.
@@ 7.1. Multiplicities of tilting objects
Let ${\mathcal{F}}\in{\mathscr{P}}({\operatorname{Hk}}_{{{\mathcal{I}}{\mathcal{W}}}})$ be a tilting object. Recall that the multiplicity of the standard (resp. costandard) objects $\Delta^{{\mathcal{I}}{\mathcal{W}}}_{\nu}$ (resp. $\nabla^{{\mathcal{I}}{\mathcal{W}}}_{\nu}$) in ${\mathcal{F}}$ is well-defined and we denote it by $(\mathcal{F}:\Delta^{{\mathcal{I}}{\mathcal{W}}}_{\nu})$(resp. $({\mathcal{F}}:\nabla^{{\mathcal{I}}{\mathcal{W}}}_{\nu}$). It follows by orthogonality of $\Delta_{\nu}^{{\mathcal{I}}{\mathcal{W}}}$ and $\nabla_{\nu}^{{\mathcal{I}}{\mathcal{W}}}$ that
$(\mathcal{F}:\Delta^{{\mathcal{I}}{\mathcal{W}}}_{\nu})\simeq\dim\mathrm{Hom}({\mathcal{F}},\nabla^{{\mathcal{I}}{\mathcal{W}}}_{\nu}),$ (7.2)
$(\mathcal{F}:\nabla^{{\mathcal{I}}{\mathcal{W}}}_{\nu})\simeq\dim\mathrm{Hom}(\Delta^{{\mathcal{I}}{\mathcal{W}}}_{\nu},{\mathcal{F}}).$ (7.3)
@@
$\sum_{i\geq 0}(-1)^{i}({\mathscr{Z}}_{{\mathcal{I}}{\mathcal{W}}}(V)[-i]:\nabla_{\nu}^{{\mathcal{I}}{\mathcal{W}}})=\dim(V(\nu))$ (7.4)
$\sum_{i\leq 0}(-1)^{i}({\mathscr{Z}}_{{\mathcal{I}}{\mathcal{W}}}(V)[i]:\Delta_{\nu}^{{\mathcal{I}}{\mathcal{W}}})=\dim(V(\nu))$ (7.5)
@@
The rest of this section is devoted to the proof of the following theorem:
Theorem 7.4.
If $V$ is a direct summand of a tensor product of minuscule representations, then ${\mathscr{Z}}_{{\mathcal{I}}{\mathcal{W}}}(V)$ is tilting.
In particular, we deduce that every ${\mathscr{Z}}_{{\mathcal{I}}{\mathcal{W}}}(V)$ is tilting if every simple normal quotient of $G$ is isomorphic to $\mathrm{PGL}_{n}$ for some (not fixed) positive integer $n$, because every weight is a sum of minuscule ones.
First, we handle minuscule representations, and for that we require the next lemma.
@@
$\mathrm{Ext}^{n}(\nabla_{\nu}^{{\mathcal{I}}{\mathcal{W}}},{\mathscr{Z}}_{{\mathcal{I}}{\mathcal{W}}}(V))\simeq\mathrm{Ext}^{n}(\nabla_{x(\nu)}^{{\mathcal{I}}{\mathcal{W}}},{\mathscr{Z}}_{{\mathcal{I}}{\mathcal{W}}}(V)).$
@@
by Proposition 6.6 and Lemma 3.2. Then

### v4

@@ 4.5. Highest weight arrows
Let $\mu\in\bar{{\mathbb{X}}}_{\bullet}$ be a dominant coweight with respect to $B$. For a $\hat{G}_{\Lambda}$-representation $V$ with a single highest weight $\mu$, we see that ${\mathscr{Z}}(V)$ is supported on the $\mu$-admissible locus ${\mathcal{A}}_{{\mathcal{I}},\mu}$, cf. [AGLR22, Theorem 6.16], which equals the union of the ${\mathcal{I}}(O)$-orbits of the translations $t_{\bar{\nu}}$ associated with weights $\nu$ of $V$. We are going to define a canonical map
@@
But the restriction of ${\mathscr{Z}}(V)$ to the $I$-orbit ${\mathrm{Fl}}_{{\mathcal{I}},\bar{\mu}}$ is isomorphic to the local system with value $R\Gamma({\mathrm{Fl}}_{{\mathcal{I}},\bar{\mu}},{\mathscr{Z}}(V))$. On the other hand, we know by Lemma 2.12 that ${\mathrm{Fl}}_{{\mathcal{I}},\bar{\mu}}$ coincides with the intersection ${\mathcal{A}}_{{\mathcal{I}},\mu}\cap{\mathcal{S}}_{t_{\bar{\mu}}}={\mathrm{Fl}}_{{\mathcal{I}},\leq\bar{\mu}}\cap{\mathcal{S}}_{t_{\bar{\mu}}}$. Therefore, Corollary 3.24 tells us that $j^{*}_{\bar{\mu}}{\mathscr{Z}}(V)\simeq\operatorname{Grad}_{\bar{\mu}}{\mathscr{Z}}(V)[\langle 2\rho,\mu\rangle]$ in natural fashion. In particular, we get $Rj_{\bar{\mu},*}j^{*}_{\bar{\mu}}{\mathscr{Z}}(V)\simeq\operatorname{gr}_{\bar{\mu}}{\mathscr{Z}}(V)$ and we obtain the desired highest weight arrow.
@@
By Corollary 4.22 we know that the composition $\operatorname{Grad}\circ{\mathscr{Z}}$ identifies with the restriction functor from $\mathrm{Rep}_{\Lambda}(\hat{G})$ to $\mathrm{Rep}_{\Lambda}(\hat{T}^{I})$ as a tensor functor, so it is symmetric monoidal. Indeed, the monoidal structure of $\operatorname{Sat}$ comes from the monoidality of constant terms in the generic fiber ${\operatorname{Gr}}_{G}$ which is compatible with the one in the special fiber ${\mathrm{Fl}}_{{\mathcal{I}}}$, which was used in Corollary 3.28. Finally, we just have to remark that the adjunction unit is naturally symmetric monoidal as are the isomorphisms $Rj_{\bar{\mu},*}j^{*}_{\bar{\mu}}{\mathscr{Z}}(V)\simeq\operatorname{gr}_{\bar{\mu}}{\mathscr{Z}}(V)$. ∎
@@
$\mathfrak{f}_{V}\circ\mathbf{n}_{V}=0.$
@@ 4.6. Mixed variant
In this subsection, we are going to upgrade our previous work to the setting of mixed sheaves. We consider a $p$-adic field $F$ with ring of integers $O$, a finite residue field $k$ of cardinality $q$, and an absolute Galois group $\Gamma$. We continue to fix a quasi-split and residually split $F$-group $G$ with an Iwahori $O$-model ${\mathcal{I}}$. In this subsection, we assume furthermore that ${\Lambda}$ is an algebraic extension of ${\mathbb{Q}}_{\ell}$ and contains a preferred choice of square-root $\sqrt{q}$.
@@
Again, we can define the mixed standard functor from ${\mathscr{D}}_{\mathrm{\acute{e}t}}([\Gamma\backslash S_{\bar{k}}])$ towards ${\mathscr{D}}_{\mathrm{\acute{e}t}}([\Gamma\backslash{\operatorname{Hk}}_{{\mathcal{I}},\bar{k}}])$
@@
both of which preserve mixed perverse sheaves by Weil II. Lemma 3.2 and Lemma 3.3 generalize to the current setting, so that we can define the mixed Wakimoto functor ${\mathscr{I}}_{\bar{\nu}}^{\mathrm{mix}}\colon{\mathscr{D}}_{\mathrm{\acute{e}t}}([\Gamma\backslash S_{\bar{k}}])\to{\mathscr{D}}_{\mathrm{\acute{e}t}}([\Gamma\backslash{\operatorname{Hk}}_{{\mathcal{I}},\bar{k}}])$ mapping a weighted complex $M$ of ${\Lambda}$-modules to the object representing
@@
where $\bar{\nu}_{1},\bar{\nu}_{2}\in\bar{\mathbb{X}}_{\bullet}^{+}$ run over all those elements such that $\bar{\nu}=\bar{\nu}_{1}-\bar{\nu}_{2}$. Again this sends a mixed weighted ${\Lambda}$-module to a mixed perverse sheaf, since the mixedness property is preserved under derived pushforward and pullback, whereas perversity was already verified in Lemma 3.6. We define the full subcategory
$\mathrm{Wak}^{\mathrm{mix}}\subset{\mathscr{D}}_{\mathrm{\acute{e}t}}([\Gamma\backslash{\operatorname{Hk}}_{{\mathcal{I}},\bar{k}}])$ (4.33)
of mixed Wakimoto complexes, as the span under cones and shifts of the essential image of ${\mathscr{I}}^{\mathrm{mix}}_{\bar{\nu}}$ for all $\bar{\nu}\in\bar{\mathbb{X}}_{\bullet}$. We also refer the reader to [CvdHS24, Definition 5.7] in the motivic equicharacteristic setting, which discusses weights in greater detail. Similarly, we can generalize Proposition 3.23 to the mixed setting, in such a way that it allows us to determine the Wakimoto grading of such an object. The full subcategory ${\mathscr{P}}(\mathrm{Wak}^{\mathrm{mix}})$ consists of the objects in $\mathrm{Wak}^{\mathrm{mix}}$ whose gradeds are all perverse.
@@
The mixed central functor ${\mathscr{Z}}^{\mathrm{mix}}$ lands in ${\mathscr{P}}(\mathrm{Wak}^{\mathrm{mix}})$. Concretely, the Wakimoto gradeds of ${\mathscr{Z}}^{\mathrm{mix}}(V)$ are canonically isomorphic to ${\mathscr{I}}_{\bar{\nu}}^{\mathrm{mix}}(V(w_{0}\bar{\nu}))$$\bar{\nu}\in\bar{\mathbb{X}}_{\bullet}$.
@@ 5.1. Generalities on coherent sheaves
Let $R$, resp. $A$ be the ring of global sections on $Y$ resp. $H$. The category of finitely generated $R$-modules (which is equivalent to $\mathrm{Coh}(Y)$) has enough projectives by considering the collection of free $R$-modules. Notice that the functor of taking $H$-invariants on $R$-modules is exact by assumption on $H$. We can deduce that the $H$-equivariant free $R$-module $V\otimes M$ with $V$ being a finite dimensional representation of $H$ is projective in $\mathrm{Coh}(X)$. This collection of projectives is enough, as each coherent sheaf on $X$ is surjected upon by the $n$-fold sum of the regular representation $A\otimes R$ for $n\gg 0$, and we can find a finite representation $V\subset A$ which completes the job, by finiteness of the underlying $R$-module of the initial coherent sheaf on $X$. ∎
@@ 6. Iwahori–Whittaker averaging
Our next task is to choose a Whittaker datum. Assume $\Lambda$ is an algebraic field extension of ${\mathbb{Q}}_{\ell}(\zeta)$ where $\zeta\in\bar{\mathbb{Q}}_{\ell}$ is a choice of a primitive $p$-th root of unity. We get the Artin–Schreier étale sheaf ${\mathcal{L}}_{\mathrm{AS}}$ on ${\mathbb{G}}_{a,k}$: this is the rank $1$ direct summand of the pushforward $\pi_{*}{\Lambda}$ of the constant sheaf along the Artin–Schreier cover $\pi\colon{\mathbb{G}}_{a,k}\to{\mathbb{G}}_{a,k}$ arising as the $\zeta$-eigenspace for the Galois action of ${\mathbb{Z}}/p{\mathbb{Z}}$. It is a character sheaf in the sense of Lusztig–Yun [LY20], i.e., we have isomorphisms $m^{*}{\mathcal{L}}_{\mathrm{AS}}\simeq{\mathcal{L}}_{\mathrm{AS}}\boxtimes{\mathcal{L}}_{\mathrm{AS}}$ and $e^{*}{\mathcal{L}}_{\mathrm{AS}}\cong{\Lambda}$ with respect to the multiplication $m$ and unit $e$ of ${\mathbb{G}}_{a,k}$, that satisfy associativity constraints (this is equivalent to the corresponding $\infty$-enhancement, because ${\mathscr{P}}({\mathbb{G}}_{a,k})$ is an abelian category). Besides, it satisfies the following cohomological vanishing
@@
The orbit ${\mathrm{Fl}}^{\mathrm{op}}_{{\mathcal{I}},w}$ carries an Iwahori–Whittaker local system if and only if $w$ has minimal length in its $W_{\mathrm{fin}}$-left coset, i.e., $\ell(w_{\mathrm{fin}}w)\geq\ell(w)$ for all $w_{\mathrm{fin}}\in W_{\mathrm{fin}}$. If the latter condition holds, then the rank $1$ Iwahori–Whittaker local system on ${\mathrm{Fl}}^{\mathrm{op}}_{{\mathcal{I}},w}$ is unique up to isomorphism.
@@
and claim that its composition with the quotient map ${\mathscr{P}}({\operatorname{Hk}}_{{\mathcal{I}}})\to{\mathscr{P}}_{\mathrm{as}}({\operatorname{Hk}}_{{\mathcal{I}}})$, which we denote by $\mathrm{ind}_{{\mathcal{I}}{\mathcal{W}}}^{\mathrm{as}}$, yields a right inverse to $\mathrm{av}^{\mathrm{as}}_{{\mathcal{I}}{\mathcal{W}}}$. First notice that $\mathrm{ind}_{{\mathcal{I}}{\mathcal{W}}}(\Xi)$ is an extension of negative shifts of $\mathrm{IC}_{w}$ for $w\in W_{\mathrm{fin}}$, with the local system $\mathrm{IC}_{1}$ appearing with multiplicity $1$ (see [AR, Lemma 6.4.8]). If $w$ is non-trivial and $\mathcal{F}\in{\mathcal{P}}({\operatorname{Hk}}_{\mathcal{I}})$, we can show that $\mathrm{IC}_{w}[n]\ast{\mathcal{F}}$ is $L^{+}{\mathcal{P}}_{s}$-equivariant for some simple reflection $s$, hence its perverse cohomology groups die under the quotient map ${\mathscr{P}}({\operatorname{Hk}}_{{\mathcal{I}}})\to{\mathscr{P}}_{\mathrm{as}}({\operatorname{Hk}}_{{\mathcal{I}}})$. If $w=1$ and $n\neq 0$, then $\mathrm{IC}_{1}[n]\ast{\mathcal{F}}$ sits in non-zero perverse degree. In total, this yields an equivalence of functors $\mathrm{ind}_{{\mathcal{I}}{\mathcal{W}}}^{\mathrm{as}}\circ\mathrm{av}_{{\mathcal{I}}{\mathcal{W}}}^{\mathrm{as}}\simeq\mathrm{id}$. This implies that $\mathrm{av}_{{\mathcal{I}}{\mathcal{W}}}^{\mathrm{as}}$ is injective on $\mathrm{Ext}$ groups. To see that $\mathrm{av}_{{\mathcal{I}}{\mathcal{W}}}^{\mathrm{as}}$ is fully faithful, we argue by induction on the length of the objects being considered: the simple case is a consequence of Lemma 6.5, implying bijectivity of simple objects along $\mathrm{av}_{{\mathcal{I}}{\mathcal{W}}}^{\mathrm{as}}$; the induction step follows from the 5-lemma and the injectivity on $\mathrm{Ext}$-groups. ∎
@@ 7. Tilting modules
lands on the full subcategory of tilting objects. This is related to the ’parabolic-singular’ Koszul duality phenomenon studied by Beilinson–Ginzburg–Soergel [BGS96] for finite flag varieties and Bezrukavnikov–Yun [BY13] for Kac–Moody flag varieties. We continue to assume that $\Lambda$ is an algebraic field extension of $\mathbb{Q}_{\ell}(\zeta_{\ell})$ for a fixed primitive $p$-th root of unity $\zeta\in\bar{\mathbb{Q}_{\ell}}$.
@@ 7.1. Multiplicities of tilting objects
Let ${\mathcal{F}}\in{\mathscr{P}}({\operatorname{Hk}}_{{{\mathcal{I}}{\mathcal{W}}}})$ be a tilting object. Recall that the multiplicity of the standard (resp. costandard) objects $\Delta^{{\mathcal{I}}{\mathcal{W}}}_{\nu}$ (resp. $\nabla^{{\mathcal{I}}{\mathcal{W}}}_{\nu}$) in ${\mathcal{F}}$ is well-defined and we denote it by $(\mathcal{F}:\Delta^{{\mathcal{I}}{\mathcal{W}}}_{\nu})$(resp. $({\mathcal{F}}:\nabla^{{\mathcal{I}}{\mathcal{W}}}_{\nu}$)). It follows by orthogonality of $\Delta_{\nu}^{{\mathcal{I}}{\mathcal{W}}}$ and $\nabla_{\nu}^{{\mathcal{I}}{\mathcal{W}}}$ that
$(\mathcal{F}:\Delta^{{\mathcal{I}}{\mathcal{W}}}_{\nu})=\dim\mathrm{Hom}({\mathcal{F}},\nabla^{{\mathcal{I}}{\mathcal{W}}}_{\nu}),$ (7.2)
$(\mathcal{F}:\nabla^{{\mathcal{I}}{\mathcal{W}}}_{\nu})=\dim\mathrm{Hom}(\Delta^{{\mathcal{I}}{\mathcal{W}}}_{\nu},{\mathcal{F}}).$ (7.3)
@@
$\sum_{i\geq 0}(-1)^{i}\dim({\operatorname{Hom}}_{{\mathscr{D}}_{\mathrm{\acute{e}t}}({\operatorname{Hk}}_{{\mathcal{I}}{\mathcal{W}}})}({\mathscr{Z}}_{{\mathcal{I}}{\mathcal{W}}}(V)[i],\nabla_{\nu}^{{\mathcal{I}}{\mathcal{W}}}))=\dim(V(\nu))$ (7.4)
$\sum_{i\leq 0}(-1)^{i}\dim({\operatorname{Hom}}_{{\mathscr{D}}_{\mathrm{\acute{e}t}}({\operatorname{Hk}}_{{\mathcal{I}}{\mathcal{W}}})}(\Delta_{\nu}^{{\mathcal{I}}{\mathcal{W}}},{\mathscr{Z}}_{{\mathcal{I}}{\mathcal{W}}}(V)[i])=\dim(V(\nu)),$ (7.5)
@@
The work that now follows will eventually lead to proving that $\mathscr{Z}_{\mathcal{I}\mathcal{W}}(V)$ is tilting for almost all groups $G$, which we explain below in Theorem 7.9. But first, we handle minuscule representations, and for that we require the next lemma.
@@
$\mathrm{Ext}^{n}({\mathscr{Z}}_{{\mathcal{I}}{\mathcal{W}}}(V),\nabla_{\nu}^{{\mathcal{I}}{\mathcal{W}}})\simeq\mathrm{Ext}^{n}({\mathscr{Z}}_{{\mathcal{I}}{\mathcal{W}}}(V),\nabla_{x(\nu)}^{{\mathcal{I}}{\mathcal{W}}}).$
@@
by Proposition 6.6 and Lemma 3.2. Thus
@@
Now, we need to describe when minuscule representations form a class of Karoubi generators for the symmetric monoidal category $\mathrm{Rep}(\hat{G})$. It will be enough for us to restrict to adjoint $G$. The argument below was partially suggested to us by Jeremy Taylor.
Lemma 7.7.
Assume $G$ is adjoint. The following are equivalent:
  1. (1)
    $\mathrm{Rep}(\hat{G})$ is generated by minuscule representations under sums, retracts and tensor products;
  2. (2)
    $\hat{G}$ admits a faithful minuscule representation;
  3. (3)
    Every simple adjoint factor of $G$ contains a minuscule coweight (i.e., it is not of type $E_{8}$, $F_{4}$ nor $G_{2}$).
Proof.
Recall that the generation property when including quasi-minuscules was observed by Ngô–Polo, see [NP01, Lemma 10.3], and we could read this off their proof. It is clear that (1) implies (3). Also, (2) implies (1) because we are working over a characteristic 00 field $\Lambda$ and hence we can invoke the Peter–Weyl theorem: explicitly, taking global sections along the faithful representation $\rho\colon\hat{G}\to\mathrm{GL}(V)$ over $\Lambda$, we see that the regular $\hat{G}$-representation is spanned by minuscule representations under sums, retracts and tensor products, so by semisimplicity the same holds for any representation of $G$. Finally, we show that (3) implies (2). It is enough to assume $G$ is almost simple and let $\mu$ be a minuscule coweight. The representation $\hat{G}\to\mathrm{GL}(V_{\mu})$ has finite central kernel, which is shared by any non-trivial representation whose weights differ from $\mu$ by an element of the coroot lattice. Varying the minuscule $\mu$, we get every single element in ${\mathbb{X}}_{\bullet}$ by [Bou68, Exercices 24–25, p. 225], which implies that the direct sum of all minuscule representations is faithful. ∎
Definition 7.8.
We say that $G$ has enough minuscules if its adjoint quotient satisfies the equivalent conditions of Lemma 7.7.
Now, we can prove the tilting property under a mild assumption involving exceptional groups. As explained during the introduction, this stems from the lack of the rotation $\mathbb{G}_{m,k}^{\mathrm{pf}}$-action in our $p$-adic context.
Theorem 7.9.
If $G$ has enough minuscules, then $\mathscr{Z}_{\mathcal{I}\mathcal{W}}(V)$ is tilting for all $V\in\mathrm{Rep}(\hat{G})$.
Proof.
First of all, we perform a reduction to the adjoint case. Recall that the adjoint map $M_{{\mathcal{I}},\mu}\to M_{{\mathcal{I}}_{\mathrm{ad}},\mu_{\mathrm{ad}}}$ is an isomorphism, so if $V$ is simple (which we may and do assume), we can naturally identify the central sheaf ${\mathscr{Z}}(V)$ with the central sheaf ${\mathscr{Z}}(V_{\hat{G}_{\mathrm{sc}}})$ constructed in ${\operatorname{Hk}}_{{\mathcal{I}}_{\mathrm{ad}}}$, see also [FS21, Section VI.11] to see that geometric Satake is compatible with adjoint quotients. The natural functor $\mathscr{P}(\mathrm{Hk}_{\mathcal{I}\mathcal{W}})\to\mathscr{P}(\mathrm{Hk}_{\mathcal{I}\mathcal{W}_{\mathrm{ad}}})$ also becomes an equivalence when restricted to a single connected component of the Hecke stack, identifying standard and costandard objects in the obvious way. In particular, the assertion can be read off the adjoint case.
Now, if $G$ is adjoint, we apply Proposition 7.5, Proposition 7.6 and Lemma 7.7 in combination to arrive at the desired result. ∎